3. Tigist wishes to mix two types of food C and D in such a way that the vitamin contents of the mixture contain at least 8 units of vitamin C and 11 units of vitamin D. Food C costs $ 60/kg and Food D costs $80/kg. Food C contains 3 units/kg of Vitamin A and 5 units / kg of Vitamin B while food D contains 4 units/kg of Vitamin A and 2 units/kg of vitamin B. Determine the minimum cost of the mixture by graphic model.

1
Expert's answer
2022-01-04T11:22:50-0500

minimize cost:

"Z=60x_1+80x_2"

subject to:

"3x_1+4x_2\ge 8" - for Vitamin A

"5x_1+2x_2\ge 11" - for Vitamin B

where

x1 is number of kg of food C,

x2 is number of kg of food D




for extreme points:

"5x_1+2x_2= 11"

"x_1=0\implies x_2=5.5"


"3x_1+4x_2= 8"

"x_2=0\implies x_1=2.67"


intersection:

"x_1=(11-2x_2)\/5"

"3(11-2x_2)\/5+4x_2= 8"

"x_2=7\/14=0.5,x_1=2"


Objective function values at extreme points:

"Z(0,5.5)=440"

"Z(2,0.5)=160"

"Z(2.67,0)=160"


The miniimum value of the objective function Z=160 occurs at 2 extreme points.

Hence, problem has multiple optimal solutions and min Z=160.